OpenStudy (freemap):

In the year 2000, the estimated population of Canadian geese in a city was 750. The Canadian geese population is expected to grow at a rate of 12% each year. In what year will the population first exceed 15,000?

OpenStudy (mathstudent55):

You need to start with the equation for exponential growth. Do you know it? If you don't know it, try looking it up and write it.

OpenStudy (freemap):

d/n / d/t = r

OpenStudy (mathstudent55):

\(y = P_i(1 + r)^x\) y = population at year x \(P_i\) = initial population r = rate of growth (written as a decimal) x = year

OpenStudy (mathstudent55):

You know the future population, y = 12,000 You also know r = 12%, and \(P_i \) = 750. The only unknown is x, the number of years.

OpenStudy (freemap):

12,000 =750(1+12)^x

OpenStudy (freemap):

133 years?

OpenStudy (mathstudent55):

No. It's a much smaller number. I noticed I made a mistake, though. The future population is 15,000, not 12,000 as I wrote above. Also, you need to convert 12% to a decimal. 12% = 0.12 This is the equation you need to solve. \(15,000 = 750(1.12)^x\) First, divide both sides by 750. Then use logs to find x.

OpenStudy (freemap):

18

OpenStudy (mathstudent55):

No. I don't get 18. Divide both sides by 750 to get this. Now take log of both sides.